# Efficient Classical Algorithms for Simulating Gaussian Boson Sampling on Graphs

### v1

@@ 2 Initialize $X_{0}$ as an arbitrary matching in $G$
5 Choose a uniformly random edge $e$ from $E$
@@
10Uniformly sample a perfect matching $E_{t}$ in the subgraph induced by $X_{t}$ (by running another MCMC on $G_{X_{t}}$). If $e\not\in E_{t}$, set $X_{t+1}=X_{t}$. If $e\in E_{t}$, set $X_{t+1}=X_{t}\backslash\{e\}$ with probability $\frac{1}{1+\lambda^{2}}$ and otherwise set $X_{t+1}=X_{t}$;
@@ 14 Output the vertex set in matching $X_{T}$;
$\Pr[X\to X\cup\{e\}]=\frac{\lambda^{2}}{1+\lambda^{2}}\cdot\frac{1}{|E|}.$
@@
$t_{\operatorname{mix}}(\epsilon)=O\left(\varrho\left(\ln{\epsilon^{-1}}+\ln{\pi_{\min}(M)}^{-1}\right)\right).$
@@
In the original canonical path method, if the graph has sufficient symmetry, we can directly compute the congestion of each transition instead of constructing an injective mapping to bound the congestion. We begin our analysis with the complete graph case, focusing on direct computation of the congestion for each individual transition. We can design special canonical path to ensure that for any transition originating from a matching of size $k$, only paths transitioning between smaller matchings $(|M|<k)$ and larger matchings $(|M|>k)$ will utilize this transition. This insight enables us to collectively compute the total congestion across all transitions originating from size $k$ matchings. Furthermore, we aim to design symmetric canonical paths such that the congestion values for symmetric transitions are the same in complete graphs. This symmetry enables us to compute the congestion of individual transitions by summing the contributions from all paths that pass through them. However, the original canonical path construction, which determines the order of components of $I\oplus F$ based on fixed vertex orderings, fails to preserve the necessary symmetry properties if the graph is not complete. To relax this limitation, we introduce an enhanced canonical path framework that permits multiple distinct paths between each state pair $(I,F)$. Specifically, our solution involves considering all possible permutations of the connected components, systematically constructing transformation sequences for each ordering (for more details, see Appendix A).
For complete graphs and complete bipartite graphs, this symmetric construction leverages the matching enumeration properties specific to each graph, enables precise congestion calculations and ultimately yields a polynomial mixing time bound through careful analysis of the path distribution and transition probabilities. The complete technical proof is provided in Appendix C.1 (for bipartite graphs) and Appendix C.3 (for non-bipartite graphs). For sufficiently dense graphs where the Hafnian of each subgraph differs from the complete graph case by at most a polynomial factor, our analysis naturally extends to establishing polynomial mixing time bounds, as formalized in Theorem 2 and Theorem 3. The complete technical proof is provided in Appendix C.2 (for bipartite graphs) and Appendix C.4 (for non-bipartite graphs).
@@
We briefly discuss these approximate uniform sampling methods for perfect matchings in Appendix B.
@@
V Numerical Experiments
We conduct experiments to compare our algorithms with prior approaches. All results and plots are obtained by numerical simulations on a 10-core Apple M2 Pro chip with 16 GB memory and an NVIDIA L4 chip with 24 GB GPU memory via python 3.11.0. We use ‘networkx’ library to store and manipulate graphs and ‘thewalrus’ library to calculate Hafnian from adjacency matrix. Specifically, the graph problems we aim to solve are defined as follows:
@@ V Numerical Experiments
We design various graphs and set their total number of vertices $n=256$, which is larger than the system size of the 144-mode quantum photonic device Jiǔzhāng [12] used in the GBS experiment solving graph problems. Specifically, the graphs we use are presented as follows:
@@
  • $G_{2}$ for densest $k$-subgraph: Vertex $i$ has edges to vertices $0,~{}1,~{}…,~{}n-1-i$, which means the degree of each vertex is decreasing. As a result, finding the 80-vertex densest $k$-subgraph essentially finds the induced subgraph of the first 80 vertices.
@@
The classical algorithms and their variants enhanced by Glauber dynamics are presented in detail in Appendix E. We empirically choose sufficient and appropriate mixing time and post-select the edge sets with right size from the iterative process of Glauber dynamics.
@@
2 Initialize $X_{0}$ as an arbitrary matching in $G$;
3Initialize $t\leftarrow 0$;
4while $t<T$ do
5Choose a uniformly random edge $e$ from $E$;
6if $e$ and $X_{t}$ form a new matching then
@@ 7  Set $M=X_{t}\cup\{e\}$;
8else if $e$ is in $X_{t}$ then
9  Set $M=X_{t}\backslash\{e\}$;
10else if $e$ has a common vertex $v$ with $e^{\prime}$ in $X_{t}$ and the remaining vertex $u$ is not covered by $X_{t}$ then
11  Set $M=X_{t}\cup\{e\}\backslash\{e^{\prime}\};$
12
13else
@@ 14  Set $M=X_{t}$;
15
16Set $X_{t+1}=M$ with prob. $\min\{1,\lambda^{|M|-|X|}\}$ and otherwise set $X_{t+1}=X_{t}$;
@@
V.1 Glauber Dynamics Verification and Comparison
We first verify that the Glauber dynamics can enhance random search and simulated annealing by outperforming their original algorithms as corresponding baselines. We plot our experimental results in Figure 2.
@@ V.1 Glauber Dynamics Verification and Comparison
It can be seen from Figure 2 that compared to the original random search and simulated annealing algorithms, substituting uniform update in each iteration with Algorithm 1, Algorithm 3, and the quantum-inspired classical algorithm all significantly improve their performance in terms of Hafnian and density values up to 3x.
(a) Hafnian Random Search on $G_{1}$ with $k=16$, $c=0.1$
(b) Hafnian Simulated Annealing on $G_{1}$ with $k=16$, $c=0.1$
(c) Density Random Search on $G_{2}$ with $k=80$, $c=0.4$
(d) Density Simulated Annealing on $G_{2}$ with $k=80$, $c=0.4$
@@
Figure 2: Glauber Dynamics Verification and Comparison. Iterations $=1000$, mixing time $\leq 10000$, fugacity $\lambda=c^{2}$, annealing parameter $\gamma=0.95$. (a) The three enhanced variants are on average $240-300\%$ higher than the original classical algorithm, and the upper confidential intervals are about 350% better. (b) The three enhanced variants are on average $20-50\%$ higher than the original classical algorithm, and the upper confidential intervals are about 55% better. (c) The two enhanced variants are on average $30-35\%$ higher than the original classical algorithm, and the upper confidential intervals are about 32% better. (d) The two enhanced variants are on average $10-15\%$ higher than the original classical algorithm, and the upper confidential intervals are about 12% better.
V.2 Score Advantage Comparison with GBS
@@
The score advantage is defined as the ratio of the maximum Hafnian/density value acquired by random search enhanced by the Glauber dynamics to that acquired by the original random search at the end of the stage with 100 iterations. We remark that too many iterations may result in collapse of score advantage to 1 since even the original approach can acquire the same optimal solution as the variants by the Glauber dynamics. We plot our experimental results enhanced by Algorithm 1 in Figure 3.
Ref. [12] used a 144-vertex complete graph with random complex weights, whereas our numerical experiment is conducted on the aforementioned 256-vertex Erdös-Rényi graph $G_{3}$ with $p=0.4$. Since all weights of the edges in the graph are real, we do not need the square of modulus. In addition, the output probability of a Gaussian Boson sampler is proportional to the square of the Hafnian while our approach has a linear dependence in Hafnian. These may explain the difference between our score advantage and [12]. As shown in the left side of Figure 3, our Hafnian score advantage up to 4x is lower than the GBS experiments [12], whereas in the right side of Figure 3 our density score advantage is comparable to theirs.
(a) Hafnian Score advantage on $G_{3}$ with $c=0.4$
(b) Density Score advantage on $G_{3}$ with $c=0.4$
Figure 3: Score advantage vs. click number: RS enhanced by Algorithm 1 / original RS. Iterations $=100$, mixing time $\leq 1000$, fugacity $\lambda=c^{2}$. (a) The left side reflects the Hafnian score advantage up to 4x. (b) The right side reflects the density score advantage up to 110%.
(a) Hafnian Random Search on $G_{4}$ with $k=16$, $c=0.4$
(b) Hafnian Simulated Annealing on $G_{4}$ with $k=16$, $c=0.4$
(c) Density Random Search on $G_{4}$ with $k=80$, $c=0.8$
(d) Density Simulated Annealing on $G_{4}$ with $k=80$, $c=0.8$
Figure 4: Double-loop Glauber Dynamics. Iterations $=1000$, mixing time $\leq 10000$, fugacity $\lambda=c^{2}$, annealing parameter $\gamma=0.95$. (a) The four enhanced variants are on average 12x - 15x higher than the original classical algorithm, and the upper confidential intervals are about 10x better. (b) The four enhanced variants are on average $50-100\%$ higher than the original classical algorithm, and the upper confidential intervals are about 70% better. (c) The three enhanced variants are on average 8% higher than the original classical algorithm, and the upper confidential intervals are about 9% better. (d) The three enhanced variants are on average $3\%$ higher than the original classical algorithm, and the upper confidential intervals are about 2% better.
V.3 Double-Loop Glauber Dynamics on Bipartite Graphs
@@
In this section, we further conduct experiments on a randomly connected bipartite graph $G_{4}$ to verify our theory on double-loop Glauber dynamics in Section IV. For the similar reason stated in Section V.1, we do not apply the quantum-inspired classical algorithm in [15] to the densest $k$-subgraph problem. We plot our experimental results in Figure 4.
@@
It can be seen from Figure 4 that compared to the original random search and simulated annealing algorithms, substituting the uniform update in each iteration with Algorithm 1, Algorithm 2, Algorithm 3, and the quantum-inspired classical algorithm all comparably improve their performance in terms of Hafnian and density values up to 10x.
VI Discussions
In this paper, we have proposed the double-loop Glauber dynamics to simulate Gaussian Boson Sampling on graphs and theoretically prove that it mixes in polynomial time on dense graphs. Numerically, we conduct experiments on graphs with 256 vertices, larger than the scales in former GBS experiments as well as classical simulations. In particular, we show that both the single-loop and double-loop Glauber dynamics outperform original classical algorithms for the max-Hafnian and densest $k$-subgraph problems with up to 10x improvement.
@@ VI Discussions
  • For sparse subgraphs, the ratio of the Hafnian to the corresponding complete subgraph’s Hafnian may become exponentially large, preventing direct application of the complete graph analysis. Can we apply our enhanced canonical path method to deriving mixing time upper bounds for more general graphs, such as graphs with high symmetry?
  • Our analysis of mixing time bounds becomes significantly more complicated for general non-bipartite graphs. A main bottleneck is uniform sampling of perfect matchings on such graphs, whose mixing time is an open problem and directly impacts the convergence guarantees of our inner-layer MCMC procedure.
  • Our double-loop MCMC algorithm can also be extended to weighted graphs by slightly adjusting the transition probabilities of the Markov chain (see Appendix D for details). However, the analysis of their mixing time becomes significantly more complicated and remains an open problem.
VII Acknowledgments
@@
In this Supplementary Information, we present detailed proofs of the propositions, theorems, and corollaries presented in the manuscript “Efficient Classical Algorithms for Simulating Gaussian Boson sampling on Graphs”. In Appendix A, we introduce the canonical path method widely studied in MCMC literature. In Appendix B, we introduce uniform sampling for perfect matchings. Our proofs for the polynomial mixing time of our double-loop Glauber dynamics on dense graphs are presented in Appendix C. In Appendix D, we introduce how to adapt our double-loop Glauber dynamics to weighted graphs. Finally, we present details of the algorithms in our numerical experiments with pseudocodes in Appendix E.
Appendix A Canonical Path Method
@@
$\Gamma:=\{\gamma_{IF}\mid I,F\in\Omega\}$
@@
$\varrho(\Gamma):=\max_{t=(M,M^{\prime})}\left\{\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{I,F:\gamma\text{ uses }t}\pi(I)\pi(F)|\gamma_{IF}|\right\}$ (10)
@@
Theorem 8 ([16]).
The mixing time of the lazy MC is bounded by $\tau_{\operatorname{mix}}(\varepsilon)\leq 2\varrho\left(2\ln\varepsilon^{-1}+\ln\pi_{\min}^{-1}\right)$, where $\varrho=\varrho(\Gamma)$ is the congestion defined in (10) with respect to any set of canonical paths $\Gamma$, and $\pi_{\min}=\min_{M\in\Omega}\pi(M)$ is the minimum value of the stationary distribution.
@@
Definition 1 (Multiple canonical path framework).
@@
ϱ(Γ):=assignitalic-$\varrho(\Gamma):=$ $\max_{t=(M,M^{\prime})}\bigg{\{}\frac{1}{\pi(M)P(M,M^{\prime})}\cdot\sum_{I,F\in\Omega}\sum_{i=1}^{l(I,F)}\textbf{1}_{\{\gamma_{IF}^{(i)}\text{ uses }t\}}\cdot\pi(I)\pi(F)|\gamma_{IF}|\cdot p^{(i)}_{IF}\bigg{\}}.$ (11)
Similarly, Theorem 8 remains valid for the enhanced multiple canonical path framework. In subsequent proofs in Appendix C, we will employ this enhanced canonical path framework, as its multiple-path construction enables more effective utilization of symmetries between different transitions, thereby providing greater flexibility in establishing conductance bounds and mixing time estimates.
The key step in the proof in Theorem 8 is constructing the relations between the congestion $\varrho$ as well as local and global variations:
Lemma 9.
@@
$\operatorname{Var}_{\pi}(f):=\sum_{M\in\Omega}\pi(M)\left(f(M)-E_{\pi}f\right)^{2}=\frac{1}{2}\sum_{I,F\in\Omega}\pi(I)\pi(F)\left(f(I)-f(F)\right)^{2}$
@@
$\mathcal{E}_{\pi}(f,f):=\frac{1}{2}\sum_{M,M^{\prime}\in\Omega}\pi(M)P(M,M^{\prime})\left(f(M)-f(M^{\prime})\right)^{2}.$

### v2

@@ 2 Initialize $X_{0}$ as an arbitrary matching in $G$;
5 Choose a uniformly random edge $e$ from $E$;
@@
[…]
@@ 14 Output the vertex set in matching $X_{T}$;
$\Pr[X\to X\cup\{e\}]=\frac{\lambda^{2}}{1+\lambda^{2}}\cdot\frac{1}{|E|}.$ (31)
@@
$t_{\operatorname{mix}}(\epsilon)=O\left(\varrho\left(\ln{\epsilon^{-1}}+\ln{\pi_{\min}(M)}^{-1}\right)\right).$ (37)
@@
In the original canonical path method, if the graph has sufficient symmetry, we can directly compute the congestion of each transition instead of constructing an injective mapping to bound the congestion. We begin our analysis with the complete graph case, focusing on direct computation of the congestion for each individual transition. We can design special canonical path to ensure that for any transition originating from a matching of size $k$, only paths transitioning between smaller matchings $(|M|<k)$ and larger matchings $(|M|>k)$ will utilize this transition. This insight enables us to collectively compute the total congestion across all transitions originating from size $k$ matchings. Furthermore, we aim to design symmetric canonical paths such that the congestion values for symmetric transitions are the same in complete graphs. This symmetry enables us to compute the congestion of individual transitions by summing the contributions from all paths that pass through them. However, the original canonical path construction, which determines the order of components of $I\oplus F$ based on fixed vertex orderings, fails to preserve the necessary symmetry properties if the graph is not complete. To relax this limitation, we introduce an enhanced canonical path framework that permits multiple distinct paths between each state pair $(I,F)$. Specifically, our solution involves considering all possible permutations of the connected components, systematically constructing transformation sequences for each ordering (for more details, see Supplementary Note 1).
For complete graphs and complete bipartite graphs, this symmetric construction leverages the matching enumeration properties specific to each graph, enables precise congestion calculations and ultimately yields a polynomial mixing time bound through careful analysis of the path distribution and transition probabilities. The complete technical proof is provided in Supplementary Note A (for bipartite graphs) and Supplementary Note C (for non-bipartite graphs). For sufficiently dense graphs where the Hafnian of each subgraph differs from the complete graph case by at most a polynomial factor, our analysis naturally extends to establishing polynomial mixing time bounds, as formalized in Theorem 2 and Theorem 3. The complete technical proof is provided in Supplementary Note B (for bipartite graphs) and Supplementary Note D (for non-bipartite graphs).
@@
We briefly discuss these approximate uniform sampling methods for perfect matchings in Supplementary Note 2.
@@
We note that our theoretical performance bound in Theorem 4 is stronger for bipartite graphs. This advantage stems from the fact that the subroutine of uniformly sampling perfect matchings is known to be classically more efficient for bipartite graphs [27]. For context, the BipartiteGBS technique also leverages the unique properties of bipartite graphs, though for the different goal of encoding arbitrary matrices for hardness proofs [32].
II.4 Numerical experiments
We conduct experiments to compare our algorithms with prior approaches. All results and plots are obtained by numerical simulations on a 10-core Apple M2 Pro chip with 16 GB memory and an NVIDIA L4 chip with 24 GB GPU memory via python 3.11.0. We use ‘networkx’ library [33] to store and manipulate graphs and ‘thewalrus’ library [34] to calculate Hafnian from adjacency matrix. Specifically, the graph problems we aim to solve are defined as follows:
@@ II.4 Numerical experiments
We design various unweighted graphs and set their total number of vertices $n=256$, which is larger than the system size of the 144-mode quantum photonic device Jiǔzhāng [12] used in the GBS experiment solving graph problems as well as classical simulations. Specifically, the graphs we use are presented as follows:
@@
  • $G_{2}$ for densest $k$-subgraph: Vertex $i$ has edges to vertices 0,1,,n1i0,~1,~…,~n-1-i, which means the degree of each vertex is decreasing. As a result, finding the 80-vertex densest $k$-subgraph essentially finds the induced subgraph of the first 80 vertices.
@@
  • $G_{5}$ for sparse graph experiments: A random bipartite graph with 128 vertices in each part, and there are $10\cdot 256$ edges chosen among all the $128^{2}$ possible pairs uniformly at random.
The classical algorithms and their variants enhanced by Glauber dynamics are presented in detail in Supplementary Note 5. We empirically choose sufficient and appropriate mixing time and post-select the edge sets with right size from the iterative process of Glauber dynamics.
@@
2
31em Initialize $X_{0}$ as an arbitrary matching in $G$;
4Initialize $t\leftarrow 0$;
5while $t<T$ do
6Choose a uniformly random edge $e$ from $E$;
7if $e$ and $X_{t}$ form a new matching then
@@ 8  Set $M=X_{t}\cup\{e\}$;
9else if $e$ is in $X_{t}$ then
[…]
11else if $e$ has a common vertex $v$ with $e^{\prime}$ in $X_{t}$ and the remaining vertex $u$ is not covered by $X_{t}$ then
[…]
13
14else
@@ 15  Set $M=X_{t}$;
16
17Set $X_{t+1}=M$ with prob. $\min\{1,\lambda^{|M|-|X|}\}$ and otherwise set $X_{t+1}=X_{t}$;
@@
II.4.1 Glauber dynamics verification and comparison
We first verify that the Glauber dynamics can enhance random search and simulated annealing by outperforming their original algorithms as corresponding baselines. We plot our experimental results in Figure 2. The instance $G_{1}$ serves as a test on the classic planted clique problem [35], while $G_{2}$ acts as a sanity check with a predictable density structure. The objective is to confirm that our algorithms can correctly identify optimal solutions.
@@ II.4.1 Glauber dynamics verification and comparison
[…]
@@
Figure 2: Glauber Dynamics Verification and Comparison. Iterations =1000=1000, mixing time $\leq 10000$, fugacity $\lambda=c^{2}$, annealing parameter $\gamma=0.95$. (a) Hafnian Random Search on $G_{1}$ with $k=16$, $c=0.1$. The three enhanced variants are on average $240-300\%$ higher than the original classical algorithm, and the upper confidential intervals are about 350% better. (b) Hafnian Simulated Annealing on $G_{1}$ with $k=16$, $c=0.1$. The three enhanced variants are on average $20-50\%$ higher than the original classical algorithm, and the upper confidential intervals are about 55% better. (c) Density Random Search on $G_{2}$ with $k=80$, $c=0.4$. The two enhanced variants are on average $30-35\%$ higher than the original classical algorithm, and the upper confidential intervals are about 32% better. (d) Density Simulated Annealing on $G_{2}$ with $k=80$, $c=0.4$. The two enhanced variants are on average $10-15\%$ higher than the original classical algorithm, and the upper confidential intervals are about 12% better.
II.4.2 Score advantage comparison with GBS
@@
To evaluate performance on a more challenging average-case scenario, we use the Erdös-Rényi random graph $G_{3}$. The score advantage is defined as the ratio of the maximum Hafnian/density value acquired by random search enhanced by the double-loop Glauber dynamics to that acquired by the original random search at the end of the stage with 100 iterations. We remark that too many iterations may result in collapse of score advantage to 1 since even the original approach can acquire the same optimal solution as the variants by the Glauber dynamics. We plot our experimental results enhanced by Algorithm 2 in Figure 3, whose probability distribution on unweighted graph is comparable to GBS in [12].
[…]
II.4.3 Double-loop Glauber dynamics on bipartite graphs
@@
In this section, we further conduct experiments on a randomly connected bipartite graph $G_{4}$ to verify our theory on double-loop Glauber dynamics in Section II.3. This experiment is crucial because bipartite graphs are common in applications and are structures where our uniform perfect matching sampling subroutine is known to be efficient [27]. For the similar reason stated in Section II.4.1, we do not apply the quantum-inspired classical algorithm in [17] to the densest $k$-subgraph problem. We plot our experimental results in Figure 4.
[…]
Figure 5: Sparse Graph Experiments. Iterations =1000=1000, mixing time $\leq 1000$, fugacity $\lambda=c^{2}$, annealing parameter $\gamma=0.95$. (a) Hafnian Random Search on $G_{5}$ with $k=16$, $c=0.6$. (b) Hafnian Simulated Annealing on $G_{5}$ with $k=16$, $c=0.6$. In both cases the four enhanced variants are capable of finding subgraphs with considerable Hafnian values, while the original random search and simulated annealing almost completely fail to find any perfect matching. (c) Density Random Search on $G_{5}$ with $k=80$, $c=0.8$. (d) Density Simulated Annealing on $G_{5}$ with $k=80$, $c=0.8$. Regarding the density values, the three enhanced variants are still $10\%$ higher than the performance of original algorithms with the same stopping time as in dense graphs, demonstrating the convergence of our MCMC-based algorithms on this sparse graph.
II.4.4 Sparse graph
In this work, we only provide theoretical guarantee of Algorithm 2 for dense graphs in terms of polynomial mixing time. As a complementary from the practical perspective, in this subsection we conduct experiments on a sparse random bipartite graph $G_{5}$ whose number of edges $m=O(n)$, while for $G_{1}$, $G_{2}$, $G_{3}$, $G_{4}$ the parameter $m$ scales with $n^{2}$.
@@ II.4.4 Sparse graph
It can be seen from Figure 5 that the original random search and simulated annealing algorithms almost completely fail to find any subgraph with perfect matching and the acquired Hafnian value is nearly zero. In contrast, substituting the uniform update in each iteration with Algorithm 1, Algorithm 2, and Algorithm 3, the classical sampling algorithms are still capable of acquiring considerable Hafnian values.
Regarding the density values, the three enhanced variants are still $10\%$ higher than the performance of original algorithms with the same stopping time as in dense graphs, demonstrating the convergence of our MCMC-based algorithms on this sparse graph.
III Discussion
[…]
@@ III Discussion
  • For non-dense subgraphs, the ratio of the Hafnian to the corresponding complete subgraph’s Hafnian may become exponentially large, preventing direct application of the complete graph analysis, as is discussed in Supplementary Note E. Can we apply our enhanced canonical path method to deriving mixing time upper bounds for more general graphs, such as graphs with high symmetry?
  • For the analysis of mixing time for non-dense non-bipartite graphs, another technical barrier lies in the uniform sampling of perfect matchings on such graphs, whose mixing time is an open problem and directly impacts the convergence guarantees of our inner-layer MCMC procedure.
  • Our double-loop MCMC algorithm can also be extended to non-negative weighted graphs by slightly adjusting the transition probabilities of the Markov chain (see Algorithm 7 in Supplementary Note 5). However, the analysis of their mixing time becomes significantly more complicated and remains an open problem.
Acknowledgments
@@
Data Availability
The data and source code are available on our Github page https://github.com/Qubit-Fernand/GBS-MCMC. In addition, they are also available at Zenodo repository [36].
Author contributions
T.L. conceived the project. Y.Z. provided the theoretical proof of the mixing time of MCMC-based algorithms with assistance from Z.Y. S.Z. conducted the numerical experiments with assistance from Z.W., R.Y., and Y.X. X.W. proposed the framework of the double-loop Glauber dynamics. All authors contributed to writing the paper. Y.Z., S.Z., X.W., and T.L. contributed to the review process.
Competing interests
The authors declare no competing interests.
@@
In this Supplementary Information, we present detailed proofs of the propositions, theorems, and corollaries presented in the manuscript “Efficient Classical Sampling from Gaussian Boson Sampling Distributions on Unweighted Graphs”. The methods used in our proofs are structured as follows. In Supplementary Note 1, we review the canonical path method, a well-established technique in MCMC literature. Supplementary Note 2 summarizes existing results on uniform sampling for perfect matchings. In Supplementary Note 3, we present the proof for the polynomial mixing time of our double-loop Glauber dynamics on dense graphs, which is our core contribution. In Supplementary Note 4, we describe how to adapt our double-loop Glauber dynamics to weighted graphs. Finally, we present details of the algorithms in our numerical experiments with pseudocodes in Supplementary Note 5.
Supplementary Note 1 Canonical Path Method
@@
$\Gamma:=\{\gamma_{IF}\mid I,F\in\Omega\}$ (S1)
@@
$\varrho(\Gamma):=\max_{t=(M,M^{\prime})}\left\{\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{I,F:\gamma\text{ uses }t}\pi(I)\pi(F)|\gamma_{IF}|\right\}$ (S2)
@@
Theorem S1 ([18]).
The mixing time of the lazy MC is bounded by $\tau_{\operatorname{mix}}(\varepsilon)\leq 2\varrho\left(2\ln\varepsilon^{-1}+\ln\pi_{\min}^{-1}\right)$, where $\varrho=\varrho(\Gamma)$ is the congestion defined in (S2) with respect to any set of canonical paths $\Gamma$, and $\pi_{\min}=\min_{M\in\Omega}\pi(M)$ is the minimum value of the stationary distribution.
@@
Definition S1 (Multiple canonical path framework).
@@
$\varrho(\Gamma):=$ $\max_{t=(M,M^{\prime})}\bigg\{\frac{1}{\pi(M)P(M,M^{\prime})}\cdot\sum_{I,F\in\Omega}\sum_{i=1}^{l(I,F)}\textbf{1}_{\{\gamma_{IF}^{(i)}\text{ uses }t\}}\cdot\pi(I)\pi(F)|\gamma_{IF}^{(i)}|\cdot p^{(i)}_{IF}\bigg\}$. (S3)
Similarly, Theorem S1 remains valid for the enhanced multiple canonical path framework. In subsequent proofs in Supplementary Note 3, we will employ this enhanced canonical path framework, as its multiple-path construction enables more effective utilization of symmetries between different transitions, thereby providing greater flexibility in establishing conductance bounds and mixing time estimates.
The key step in the proof in Theorem S1 is constructing the relations between the congestion $\varrho$ as well as local and global variations:
Lemma S3.
@@
$\operatorname{Var}_{\pi}(f):=\sum_{M\in\Omega}\pi(M)\left(f(M)-E_{\pi}f\right)^{2}=\frac{1}{2}\sum_{I,F\in\Omega}\pi(I)\pi(F)\left(f(I)-f(F)\right)^{2}$ (S4)
@@
$\mathcal{E}_{\pi}(f,f):=\frac{1}{2}\sum_{M,M^{\prime}\in\Omega}\pi(M)P(M,M^{\prime})\left(f(M)-f(M^{\prime})\right)^{2}.$ (S5)